OpenStudy (dmezzullo):

Use the distributive property to simplify the expression below: one over four(g − 12) one over fourg − 3 one over fourg − 12 one over fourg − 4 one over fourg +

OpenStudy (dmezzullo):

HELP

OpenStudy (dmezzullo):

ill give medals

OpenStudy (tyteen4a03):

Multiply both g and -12 by 1/4, this is what the distributive property is about.

OpenStudy (dmezzullo):

k

OpenStudy (dmezzullo):

1/4g-12?

OpenStudy (dmezzullo):

help

OpenStudy (ash2326):

\[\frac 14 ( g-12)\] According to Distributive property \[A(B+C)=A\times B +A\times C\] \[\frac 14 (g-12)\]\[=\frac 1 4\times g- \frac 14 \times 12\] Now would simplify this?

OpenStudy (dmezzullo):

huh?

OpenStudy (dmezzullo):

simplier terms plz

OpenStudy (tyteen4a03):

This is the simplest term possible. Note that how A is "distributed" to B and C.

OpenStudy (dmezzullo):

so wats the aswer?

OpenStudy (anonymous):

0.25-3

OpenStudy (dmezzullo):

?

OpenStudy (anonymous):

1/4 simplified is 0.25 so 0.25(g-12)